# SuperM2M: Supervised and Mixture-to-Mixture Co-Learning for Speech Enhancement and Robust ASR

### v2

@@ 7.7 Effects of Speaker Reinforcement
In row $8$a-$8$c of Table 6.7.2 and 7, where the Whisper ASR model is used for recognition, we apply speaker reinforcement with the energy level $\gamma$ between the enhancement output and input mixture tuned based on the set of $\{10,15,20\}$ dB. Better ASR performance is observed in the $1$-channel case but not in the $6$-channel case, possibly because the enhanced speech is already reliable in the $6$-channel case, rendering speaker reinforcement not necessary.
In Table 7.5, where the ASR system proposed in [53] is used for recognition, applying speaker reinforcement in row $2$b, $4$b and $6$b respectively outperforms $2$a, $4$a and $6$a, pushing down the WER on the real test set to $2.40$%, $1.84$% and $1.58$%.
@@ 7.8 Comparison with Using Close-Talk Mixtures for ASR
In row $7$ of Table 7.5 and row $9$ of Table 7 and 6.7.2, we provide the ASR results of using close-talk mixtures for decoding. We observe that our proposed $6$-channel system obtains ASR results comparable to using close-talk mixtures for decoding. This further indicates the effectiveness of SuperM2M and our robust ASR system.
@@ 8 Conclusion
We have proposed to adapt M2M training for neural speech enhancement, where the models can be trained on real-recorded far-field mixtures in an unsupervised way, and on real-recorded close-talk and far-field mixture pairs in a weakly-supervised way. To improve M2M training, we have proposed SuperM2M, a novel co-learning algorithm that trains neural speech enhancement models by alternating between supervised training on simulated data and M2M training on real data. Evaluation results on the challenging CHiME-4 benchmark show the effectiveness of SuperM2M for speech enhancement and robust ASR. Future research will modify and evaluate SuperM2M on conversational speech separation and recognition.

### v3

@@ 8.7 Effects of Speaker Reinforcement
In system $8$a-$8$c of Table 7.7.1 and 8, where the Whisper ASR model is used for recognition, we apply speaker reinforcement with the energy level $\gamma$ between the enhancement output and input mixture tuned based on the set of $\{10,15,20\}$ dB. Better ASR performance is observed in the $1$-channel case but not in the $6$-channel case, possibly because the enhanced speech is already reliable in the $6$-channel case, rendering speaker reinforcement not necessary.
In Table 8.5, where the ASR system proposed in [53] is used for recognition, applying speaker reinforcement in system $2$b, $4$b and $6$b respectively outperforms $2$a, $4$a and $6$a, pushing down the WER on the real test set to $2.40$%, $1.84$% and $1.58$%.
@@ 8.8 Comparison with Using Close-Talk Mixtures for ASR
In system $7$ of Table 8.5 and system $9$ of Table 7.7.1 and 8, we provide the ASR results of using close-talk mixtures for decoding. We observe that our proposed $6$-channel system obtains ASR results comparable to using close-talk mixtures for decoding. This further indicates the effectiveness of SuperM2M and our robust ASR system.
8.9 Whisper vs. ESPnet ASR Model under Same Text Norm.
In previous tables, the WER results of using the Whisper ASR model and using the ESPnet ASR model cannot be directly compared, due to the use of different text normalization. As a side result, we compare the ESPnet ASR model with the Whisper Large v2 model, by traininig and evaluating the ESPnet ASR model using the same text normalizer as Whisper. Table 6 reports the results. We observe that the ESPnet ASR model obtains clearly better WERs than Whisper on the CHiME-4 dataset.
Table 6: ASR results on CHiME-4 obtained by using Whisper Large v2 vs. ESPnet ASR model [53] under same Whisper text normalization.
Official CHiME-4 Test Utterances
Cross reference ASR model (same Whisper text norm.) Input #mics Val. WER (%$\downarrow$ Test WER (%$\downarrow$
System Approach SIMU REAL SIMU REAL
1a - Mixture Whisper Large v2 1 7.429 816 069 699 9037.4298160696999037.429\,816\,069\,699\,9037.429 816 069 699 903 4.957 444 233 794 5224.9574442337945224.957\,444\,233\,794\,5224.957 444 233 794 522 10.969 809 760 132 3410.9698097601323410.969\,809\,760\,132\,3410.969 809 760 132 34 7.693 500 853 109 9737.6935008531099737.693\,500\,853\,109\,9737.693 500 853 109 973
1b - Mixture ESPnet ASR Model [53] 1 6.937 721 845 756 6966.9377218457566966.937\,721\,845\,756\,6966.937 721 845 756 696 4.937 275 624 218 466 54.93727562421846654.937\,275\,624\,218\,466\,54.937 275 624 218 466 5 9.599 875 930 521 0929.5998759305210929.599\,875\,930\,521\,0929.599 875 930 521 092 5.423 711 286 903 469 55.42371128690346955.423\,711\,286\,903\,469\,55.423 711 286 903 469 5
2a 7c of Table 8 SuperM2M Whisper Large v2 1 7.046 627 944 498 2257.0466279444982257.046\,627\,944\,498\,2257.046 627 944 498 225 4.783 994 191 440 4424.7839941914404424.783\,994\,191\,440\,4424.783 994 191 440 442 11.342 018 196 856 90711.34201819685690711.342\,018\,196\,856\,90711.342 018 196 856 907 6.023 473 450 183 5486.0234734501835486.023\,473\,450\,183\,5486.023 473 450 183 548
2b 7c of Table 8 SuperM2M ESPnet ASR Model [53] 1 3.956 921 587 608 9063.9569215876089063.956\,921\,587\,608\,9063.956 921 587 608 906 2.295 187 769 755 152.295187769755152.295\,187\,769\,755\,152.295 187 769 755 15 7.837 055 417 700 587.837055417700587.837\,055\,417\,700\,587.837 055 417 700 58 3.650 276 614 445 9963.6502766144459963.650\,276\,614\,445\,9963.650 276 614 445 996
3a 7c of Table 7.7.1 SuperM2M Whisper Large v2 6 3.384 156 179 412 7143.3841561794127143.384\,156\,179\,412\,7143.384 156 179 412 714 3.807 833 487 959 340 53.80783348795934053.807\,833\,487\,959\,340\,53.807 833 487 959 340 5 3.773 779 983 457 4033.7737799834574033.773\,779\,983\,457\,4033.773 779 983 457 403 4.043 224 238 663 9784.0432242386639784.043\,224\,238\,663\,9784.043 224 238 663 978
3b 7c of Table 7.7.1 SuperM2M ESPnet ASR Model [53] 6 1.189 899 967 731 526 31.18989996773152631.189\,899\,967\,731\,526\,31.189 899 967 731 526 3 1.629 623 653 745 3111.6296236537453111.629\,623\,653\,745\,3111.629 623 653 745 311 1.452 646 815 550 041 41.45264681555004141.452\,646\,815\,550\,041\,41.452 646 815 550 041 4 1.902 693 759 371 283 71.90269375937128371.902\,693\,759\,371\,283\,71.902 693 759 371 283 7
@@ 9 Conclusion
We have proposed to adapt UNSSOR and M2M training for neural speech enhancement, where the models can be trained on real-recorded far-field mixtures in an unsupervised way, and on real-recorded close-talk and far-field mixture pairs in a weakly-supervised way. To improve UNSSOR and M2M training, we have proposed SuperM2M, a novel co-learning algorithm that trains neural speech enhancement models by alternating between supervised training on simulated data and UNSSOR/M2M training on real data. Evaluation results on the challenging CHiME-4 benchmark show the effectiveness of SuperM2M for speech enhancement and robust ASR. Future research will modify and evaluate SuperM2M on conversational speech separation and recognition.